OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#5778 — Switching optical loop (LQN, LZI, ASC, TRI, HEL)
Scheduled Maintenance Report for Web Cloud
Completed
Hello,

We will switch Lesquin sites (LQN, IFS), Villeneuve d'Ascq (ASC, TRI) and Hellemmes (HEL) on the new optical infrastructure on the night between Wednesday and Thursday.
Taskworks will begin Wednesday night at 0:00. The window of intervention is scheduled between 00:00 and 6:00. The cut for customers located in these sites will be limited. Unlike the migration works of the last week, we will take a softer migration generally seamless to the sites.
We will turn on the sites one by one. We will regularly update the task work on order to keep you informed about the progress of work.

Update(s):

Date: 2011-09-13 22:38:04 UTC
The operation is delayed to the next week. We will inform you later about the accurate date of migration in this worktask.
Posted Sep 12, 2011 - 21:38 UTC